What is the length of Arc AB? The diameter goes from A to B and it is 180 degrees. The radius is 14 ft.(from A to the center)
Found 2 solutions by venugopalramana, cleomenius:
Answer by venugopalramana(3286)

TOTAL PERIMETER OF CIRCLE FOR ALL 360 DEGREES=2(PI)*RADIUS=(44/7)*RADIUS
NOW AB SPANS 180 DEG....SO ITS ARC LENGTH =(44/7)*14*180/360=44 FEET

Answer by cleomenius(959)

The length should be 180/360 (2 * pi * r)
1/2 87.96 = 43.98
Cleomenius